OpenStudy (anonymous):

Find the rate of convergence: lim(h->inf) (sinh- hcosh)/h = 0 (using taylor series)

OpenStudy (abb0t):

Hint: \[\frac{ \sin(h) }{ h } = 1\]

OpenStudy (anonymous):

so that simplifies to... lim(h->0) 1-cos(h) = 0

OpenStudy (anonymous):

taylor expansion of (1-cos(x))... 1-cos(x) + sin(x)h + cos(x)h^2/2! - sin(x)h^3/3! - cos(x)h^4/4!... (?)

OpenStudy (anonymous):

which then... 1-1 + 0 + h^2/2! - 0

OpenStudy (anonymous):

so rate of convergence = O(h^2) [i think?]